# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

### v1

@@ 1 Introduction
We investigate the existence of positive entire large solutions for the radial semilinear elliptic system
@@
    (Weights) The functions $p,q:[0,\infty)\rightarrow[0,\infty)$ are continuous, not both identically zero. We define the Green-type radial potentials
@@
    and $P(\infty)=\lim_{r\rightarrow\infty}P(r)$, $Q(\infty)=\lim_{r\rightarrow\infty}Q(r)$. We assume $\min(p,q)$ is not compactly supported.
@@
    (Monotonicity) The nonlinearities $f,g:[0,\infty)\times[0,\infty)\rightarrow[0,\infty)$ are nondecreasing in each variable, satisfy $f\left(s,t\right)>0$ and $g\left(s,t\right)>0$ $\forall s,t\in\left(0,\infty\right)$, $f(0,0)=g(0,0)=0$, are continuous on $0,\infty)^{2}$ and differentiable on $(0,\infty)^{2}$.
@@
    (Boundedness and Keller–Osserman envelope) There exists a continuous, nondecreasing function $h:[0,\infty)\rightarrow[0,\infty)$ with $h\left(s\right)>0$ $\forall s\in\left(0,\infty\right)$, $h\left(0\right)=0$ and constants $\eta>0$ sufficiently small such that
    $f(s,t)+g(s,t)\geq h(s+t),\quad\forall s,t\geq\eta.$
    Moreover, the Keller–Osserman integral
    $H\left(\infty\right)=\lim_{s\rightarrow\infty}\int_{1}^{s}\frac{dt}{\sqrt{\int_{0}^{t}h(z)\,dz}}<\infty,\qquad$ (1.2)
@@
Our aim is to extend the existence results of Lair [3] originally established for power-type nonlinearities to a broad class of general nonlinearities satisfying a Keller–Osserman-type growth condition. This generalization unifies the treatment of critical and supercritical growth regimes and provides a framework that encompasses earlier results such as those of Cîrstea and Rădulescu [1], Keller [2], Osserman [5], and others.
A positive entire solution of (1.1) is a pair $(u,v)$ of $C^{2}$ radial functions such that $u(r)>0$ and $v(r)>0$ for all $r\geq 0$. We denote $\mathbb{R}^{+}=[0,\infty)$ and define the set of central values
@@
The structure of $\mathcal{G}$ plays a central role in our analysis. In particular, we establish monotonicity properties, boundedness, and closedness of certain subsets of $\mathcal{G}$, and we identify a boundary set where all corresponding solutions are large, i.e., at least one component diverges to infinity as $r\rightarrow\infty$.
Our approach adapts the barrier method and radial integration techniques of Lair [4, 3], replacing explicit power-law growth with a general monotone envelope function $h$ (and implicitly $f+g$) satisfying the Keller–Osserman condition. This allows us to recover, as special cases, known results for power-type nonlinearities [3] and to connect with complementary bounded-solution results such as those in [8, 6, 7].
Paper structure. The paper is organized as follows: Section 2 presents our main results. Section 3 contains the detailed proofs of these results. The paper ends with Section 4, where we present our concluding remarks.
@@ 2 Main Results
Moreover, if
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)}$
and $H^{-1}$ denotes its inverse, then for all sufficiently large $r$,
@@
then
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$
@@ 3 Proof of the Theorems
We begin by recalling that any radial $C^{2}$ solution $(u,v)$ of (1.1) satisfies the equivalent system of Volterra-type integral equations
@@
valid for all $r\geq 0$. Here $\alpha_{0}=u(0)$ and $\beta_{0}=v(0)$ are the central values. It is clear that for $\alpha_{0},\beta_{0}\in(0,\infty)$, any solution of (3.1)–(3.2) is also a solution of (1.1), although it may not necessarily satisfy the large-solution condition (2.1).
Proof of Theorem 2.1
@@
The argument follows the monotone iteration scheme of [3], adapted to the present setting where $f$ and $g$ satisfy the general hypotheses (H2)–(H3) rather than being pure power functions.
Step 1: Construction of the iterative sequences. Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
$u_{0}(r)$ $\equiv\alpha_{0}>0,$ $v_{0}(r)$ $\equiv\beta_{0}>0,$
@@
$v_{k}(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big{(}u_{k-1}(s),v_{k-1}(s)\big{)}\,ds\,dt,$
for $k\geq 1$.
Step 2: Monotonicity. By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
@@
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing.
Step 3: Uniform upper bound via a scalar supersolution. The Keller–Osserman condition (H2) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
$\Delta w=(p+q)\big{(}f(w,w)+g(w,w)\big{)}\quad\text{in }[0,\infty),\quad w(0)>\alpha_{0}+\beta_{0},\quad\lim_{r\to\infty}w(r)=\infty.$ (3.3)
Such a $w$ can be constructed by solving the associated scalar radial problem with initial value $w(0)$ large enough and using the Keller–Osserman integral to guarantee blow-up at infinity.
@@
This provides a uniform bound on the iterates on every compact interval.
Step 4: Convergence. Since the sequences are monotone increasing and uniformly bounded above on compact intervals, they converge pointwise to limits
$u(r):=\lim_{k\rightarrow\infty}u_{k}(r),\quad v(r):=\lim_{k\rightarrow\infty}v_{k}(r),\quad r\geq 0.$
By monotone convergence, the limits $(u,v)$ satisfy (3.1)–(3.2) for all $r\geq 0$.
Step 5: Conclusion. The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty.
Step 6: Downward closure property. If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments, the iteration converges to a positive entire solution, hence $(\alpha,\beta)\in\mathcal{G}$.
This completes the proof of Theorem 2.1. ∎
@@
We divide the proof into two parts: boundedness and closedness of $\mathcal{G}_{\delta}$.
Boundedness. Suppose, for contradiction, that $\mathcal{G}_{\delta}$ is unbounded in $(0,\infty)\times(0,\infty)$. From Theorem 2.1, we know that $\mathcal{G}$ is downward closed in each coordinate: if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
@@
Without loss of generality, assume that
@@
Step 1: Choice of a comparison function. Let $\widetilde{p}(r):=\min\{p(r),q(r)\}$. By assumption (H1), $\widetilde{p}$ is continuous, radial, and not compactly supported. Thus, there exists $R>0$ such that $\widetilde{p}(R)>0$.
From [4], there exists a large, nonnegative radial solution $w$ of
@@
Here “large” means that $w$ blows up at the finite radius $R$.
Step 2: Construction of a solution in $\mathcal{G}_{\delta}$. Since $\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}$, pick $\alpha_{0}>\delta$ and consider the initial data
@@
By Theorem 2.1, there exists a positive entire solution $(u,v)$ of (3.1)–(3.2) with these initial values, i.e.,
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big{(}u(s),v(s)\big{)}\,ds\,dt,\quad 0\leq r<\infty,$
$v(r)$ $=\delta+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big{(}u(s),v(s)\big{)}\,ds\,dt,\quad 0\leq r<\infty.$
@@
Step 4: Comparison with $w$. The function $w$ satisfies $\Delta w=\widetilde{p}(r)\,h(w)$ in $[0,R)$, $w(0)=\delta$, and blows up at $r=R$. By the strong maximum principle (applied to u+vwu+v-w), if $u(0)+v(0)>w(0)$, then $u+v>w$ in a neighborhood of $r=0$. However, since $w$ blows up at $R$ while $u+v$ is $C^{2}$ on $[0,R)$, the comparison principle forces $u+v\leq w$ in $[0,R)$. This is a contradiction because
$u(0)+v(0)=\alpha_{0}+\delta>\delta=w(0).$
Therefore, our assumption that $\mathcal{G}_{\delta}$ is unbounded must be false, and $\mathcal{G}_{\delta}$ is bounded.
Closedness. Let $\{(\alpha_{m},\beta_{m})\}\subset\mathcal{G}_{\delta}$ be a sequence converging to $(\alpha,\beta)\in\overline{\mathcal{G}_{\delta}}$. For each $m$, Theorem 2.1 guarantees the existence of a positive entire solution $(u_{m},v_{m})$ of (3.1)–(3.2) with
@@
From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\alpha_{m}+\beta_{m}$, we deduce that the family $\{(u_{m},v_{m})\}$ is uniformly bounded on compact intervals and equicontinuous. By the Arzelá –Ascoli theorem, there exists a subsequence converging uniformly on compact sets to a limit $(u,v)$.
Passing to the limit in (3.1)–(3.2) shows that $(u,v)$ is a positive entire solution with
@@
Since $\min\{\alpha,\beta\}\geq\delta$ by construction, we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$. Thus $\mathcal{G}_{\delta}$ is closed.
This completes the proof of Theorem 2.2. ∎
@@
Standing assumptions and notation. We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
@@
Define the Keller–Osserman transform
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)},\quad s>0.$
By construction, $H$ is strictly decreasing,
$H^{\prime}(s)=\frac{-1}{f(s,s)+g\left(s,s\right)}<0,$
and $H^{\prime\prime}(s)\geq 0$ since $s\rightarrow f(s,s)+g\left(s,s\right)$ is increasing.
Step 1: Perturbation of initial data and blow-up for approximants. From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
Take $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. For each $k\geq 1$, consider the perturbed initial data
@@
If this pair belonged to $\mathcal{G}$, then by downward closure and the definition of $\mathcal{G}_{\delta}$, $(\alpha,\beta)$ would lie in the interior of $\mathcal{G}_{\delta}$, contradicting $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. Hence $(\alpha+\frac{1}{k},\beta+\frac{1}{k})\notin\mathcal{G}$.
@@
Step 2: Reduction to a scalar inequality. Let
@@
where $(u,v)$ is the limit solution corresponding to $(\alpha,\beta)$. Since $u_{k}\leq W_{k}$ and $v_{k}\leq W_{k}$, and $f,g$ are nondecreasing,
@@
$\leq(p(r)+q(r))\,\left(f(W_{k},W_{k})+g\left(W_{k},W_{k}\right)\right),$ (3.4)
where we used (H2) in the last step.
Step 3: Applying the Keller–Osserman transform. Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
@@
Since $H^{\prime}(s)=-1/\left(f(s,s)+g\left(s,s\right)\right)$, inequality (3.4) yields
@@
Step 4: First radial integration. In radial coordinates, the above inequality becomes
@@
Step 5: Second radial integration and lower bound. Integrating from $r$ to $R_{k}$ gives
@@
Letting $r\rightarrow R_{k}^{-}$, we have $W_{k}(R_{k})\rightarrow\infty$ and hence $H(W_{k}(R_{k}))\rightarrow 0$. Thus
@@
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2).
Step 6: Divergence of at least one component. Proceeding as in Step 1 of the proof of Theorem 2.2, we have already shown that
@@
If both $u$ and $v$ were bounded on $0,\infty)$, say $u(r)\leq M_{1}$ and $v(r)\leq M_{2}$ for all $r\geq 0$, then
@@
contradicting $W(r)\rightarrow+\infty$. Hence at least one of $u,v$ must diverge:
@@
Both components diverge under the supplementary Keller–Osserman condition. Assume now the additional Keller–Osserman condition at infinity (2.3). We prove
Clearly, if $u$ were bounded, then $v=W-u\rightarrow+\infty$. Thus it suffices to show
@@
Proof. Assume by contradiction that $u$ is bounded: there exists $U>0$ with $0<u(r)\leq U$ for all $r\geq 0$. Since $v(r)\rightarrow+\infty$ while $u$ is bounded, there exists $R_{0}>0$ such that
$u(r)\leq U<v(r)\qquad\forall r\geq R_{0}.$
By the monotonicity of $f$ in each variable,
$f\bigl{(}u(r),v(r)\bigr{)}>f\bigl{(}u(r),u(r)\bigr{)}=:\ell_{f}\bigl{(}u(r)\bigr{)}\qquad\forall r\geq R_{0},$
where $\ell_{f}(s):=f(s,s)$ is continuous, nondecreasing, $\ell_{f}(0)=0$, and satisfies the Keller–Osserman condition
@@
$\Delta u(r)>p(r)\,\ell_{f}\bigl{(}u(r)\bigr{)}.$ (3.5)
We now carry out a Keller–Osserman barrier construction with smaller initial data. Fix $R_{0}>0$ as above and consider the radial Cauchy problem
$y^{\prime\prime}(r)+\dfrac{n-1}{r}y^{\prime}(r)$ $=$ $p(r)\,\ell_{f}\!\big{(}y(r)\big{)},\text{ in }\left[R_{0},R_{a}\right)$ (3.6)

### v2

@@ 1 Introduction
The search for positive entire large solutions to nonlinear elliptic systems is a classical theme in analysis, with deep connections to geometry, physics, and population models. In this work, we focus on the radial semilinear system
@@
    Weights: $p,q\geq 0$ are continuous, not both trivial, $\min(p,q)$ not compactly supported, with finite Green-type potentials
@@
    and
    $P(\infty)=\lim_{r\rightarrow\infty}P(r),Q(\infty)=\lim_{r\rightarrow\infty}Q(r).$
@@
    Monotonicity: $f,g$ are continuous on $0,\infty)^{2}$, $C^{1}$ on $(0,\infty)^{2}$, positive on $(0,\infty)^{2}$, vanish at $(0,0)$, and are nondecreasing in each variable.
@@
    Keller–Osserman envelope: There exists a continuous, nondecreasing $h\geq 0$, with $h(s)>0$ for $s>0$, such that for some small $\eta>0$,
    $f(s,t)+g(s,t)\geq h(s+t),\quad\forall s,t\geq\eta,$
    and the Keller–Osserman integral
    $H(\infty)=\lim_{s\rightarrow\infty}\int_{1}^{s}\frac{dt}{\sqrt{\int_{0}^{t}h(z)\,dz}}<\infty$
@@
Our aim is to generalize the existence theory of Lair [3] "originally for power-type nonlinearities" to a broad class of nonlinearities governed by the Keller–Osserman growth condition. This unifies the treatment of critical and supercritical regimes, and recovers as special cases the results of Keller [2], Osserman [5], Cîrstea and Rădulescu [1], among others.
We call $(u,v)$ a positive entire solution if it is $C^{2}$, radial, and strictly positive on $0,\infty)$. The set of central values
@@
encodes the initial data leading to such solutions. We show that $\mathcal{G}$ has a monotone structure, certain subsets are bounded and closed, and that on its boundary, all solutions are large at least one component diverges as $r\rightarrow\infty$.
Our method adapts Lair’s barrier and radial integration techniques [4, 3], replacing explicit power laws with a general monotone envelope $h$ satisfying the Keller–Osserman condition. This approach not only recovers known power-type results but also links naturally to complementary bounded-solution theories [8, 6, 7].
Structure of the paper. Section 2 states the main results. Section 3 contains the proofs, built on monotone iteration, Keller–Osserman estimates, and barrier arguments. Section 4 concludes with remarks and perspectives.
@@ 2 Main Results
The geometric relationship between these sets is depicted in Figure 1, where $\mathcal{G}$ is shown in green, $\mathcal{G}_{\delta}$ in yellow, and the interior boundary arc $\mathcal{E}(\mathcal{G}_{\delta})$ in blue.
[…]
Figure 1: Existence region in the $(\alpha,\beta)$-plane: $G$ (green), $G_{\delta}$ (yellow), and $E(G_{\delta})$ (blue).
@@
Moreover, if $H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)}$ and $H^{-1}$ denotes its inverse, then for all sufficiently large $r$,
@@
then $\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty$.
@@
Remark 2.6.
The sets $\mathcal{G}$, $\mathcal{G}_{\delta}$, and $\mathcal{E}(\mathcal{G}_{\delta})$ can be thought of as nested regions: $\mathcal{G}$ contains all starting values that lead to positive solutions, $\mathcal{G}_{\delta}$ is a smaller core within it, and $\mathcal{E}(\mathcal{G}_{\delta})$ marks the area where the solutions grow without bound.
@@ 3 Proof of the Theorems
We recall that any radial $C^{2}$ solution $(u,v)$ of (1.1) satisfies the equivalent Volterra-type integral system
@@
for all $r\geq 0$, where $\alpha_{0}=u(0)$ and $\beta_{0}=v(0)$ are the central values. For $\alpha_{0},\beta_{0}\in(0,\infty)$, any solution of (3.1)–(3.2) is also a solution of (1.1), though it may not satisfy the large-solution condition (2.1).
3.1 Proof of Theorem 2.1
The proof of Theorem 2.1 is organized into six logically steps. We adapt the monotone iteration method of [3] to the present setting, where $f$ and $g$ satisfy the general hypotheses (H2)–(H3).
Step 1: Iterative scheme.
@@
Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
$u_{0}(r)\equiv\alpha_{0},\quad v_{0}(r)\equiv\beta_{0},$
and for $k\geq 1$ set
@@
$v_{k}(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big(u_{k-1}(s),v_{k-1}(s)\big)\,ds\,dt.$
Step 2: Monotonicity.
Proof.
By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
@@
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing. ∎
Step 3: Uniform upper bound via a scalar supersolution.
Proof.
Construction of $w$. The Keller–Osserman condition (H3) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
$\Delta w=(p+q)\big(f(w,w)+g(w,w)\big),\quad w(0)>\alpha_{0}+\beta_{0},\quad\lim_{r\rightarrow\infty}w(r)=\infty.$ (3.3)
To construct $w$, consider the scalar radial ODE
$W^{\prime\prime}(r)+\frac{n-1}{r}W^{\prime}(r)=(p(r)+q(r))\,S(W(r)),$
where $S(s):=f(s,s)+g(s,s)$. By (H2)–(H3), $H$ is continuous, nondecreasing, positive for $s>0$, and satisfies the Keller–Osserman integral condition
$\int_{1}^{\infty}\frac{ds}{\sqrt{\int_{0}^{s}S(\tau)\,d\tau}}<\infty.$
Choosing $W(0)$ sufficiently large, standard ODE theory and the above integral condition guarantee that $W(r)\rightarrow\infty$ as $r\rightarrow\infty$. Setting $w:=W$ yields (3.3).
Significance of $w$. The function $w$ is a scalar supersolution for the system: it bounds the sum $u+v$ from above and grows unboundedly at infinity. This is crucial because:
  • It provides a uniform a priori bound on the iterates $u_{k}+v_{k}$, independent of $k$, on every compact interval.
  • It allows the use of comparison principles to control the nonlinear coupling in the system by a single scalar equation.
  • It ensures that the iteration cannot “blow up” prematurely on finite intervals.
@@
Step 4: Convergence of the iteration.
Proof.
For each fixed $r\geq 0$, the sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are monotone increasing and bounded above by $w(r)$. Thus, the pointwise limits
$u(r):=\lim_{k\rightarrow\infty}u_{k}(r),\quad v(r):=\lim_{k\rightarrow\infty}v_{k}(r)$
exist for all $r\geq 0$.
Monotone Convergence Theorem (MCT). The MCT states: If $\{s_{k}\}$ is a sequence of nonnegative measurable functions such that $s_{k}(x)\uparrow s(x)$ pointwise, then
$\lim_{k\rightarrow\infty}\int s_{k}=\int\lim_{k\rightarrow\infty}s_{k}.$
In our setting, $s_{k}$ corresponds to the integrands in (3.1)–(3.2). Since $u_{k},v_{k}$ increase pointwise and $p,q,f,g$ are nonnegative, the integrands are nonnegative and increasing in $k$. Thus, MCT justifies passing the limit $k\rightarrow\infty$ inside the integrals, showing that $(u,v)$ satisfies (3.1)–(3.2). ∎
Step 5: Conclusion.
Proof.
The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty. ∎
Step 6: Downward closure property.
Proof.
If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments, the iteration converges to a positive entire solution, hence $(\alpha,\beta)\in\mathcal{G}$. ∎
@@
We divide the proof into two logically independent parts: (i) boundedness of $\mathcal{G}_{\delta}$, and (ii) closedness of $\mathcal{G}_{\delta}$.
Part I: Boundedness of $\mathcal{G}_{\delta}$
@@
Step 0: Contradiction setup. Suppose, for contradiction, that $\mathcal{G}_{\delta}$ is unbounded in $(0,\infty)\times(0,\infty)$. From Theorem 2.1, we know that $\mathcal{G}$ is downward closed in each coordinate: if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
@@
Without loss of generality, assume
@@
Step 1: Choice of a scalar comparison function. Define
$\widetilde{p}(r):=\min\{p(r),\,q(r)\}.$
By assumption (H1), $\widetilde{p}$ is continuous, radial, and not compactly supported. Hence there exists $R>0$ such that $\widetilde{p}(R)>0$. From [4] (and the Keller–Osserman framework), there exists a large nonnegative radial solution $w$ of
@@
Here large means that $w$ blows up at the finite radius $R$.
Significance of $w$:
The function $w$ is a scalar supersolution for the coupled system: it satisfies the same type of equation as $u+v$ but with the minimal weight $\widetilde{p}$ and the scalar nonlinearity $h$. It will serve as a barrier function in the comparison argument: if $u+v$ starts above $w$ at $r=0$, the maximum principle forces $u+v$ to stay above $w$ until blow-up, which is impossible if $u+v$ is smooth on $[0,R)$.
Step 2: Construction of a candidate solution in $\mathcal{G}_{\delta}$. Since $\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}$, pick $\alpha_{0}>\delta$ and consider the initial data
@@
By Theorem 2.1, there exists a positive entire solution $(u,v)$ of (3.1)–(3.2) with these initial values:
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f(u(s),v(s))\,ds\,dt,$
$v(r)$ $=\delta+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g(u(s),v(s))\,ds\,dt,$
for $0\leq r<\infty$.
@@
Step 4: Comparison with $w$ and contradiction. The function $w$ satisfies $\Delta w=\widetilde{p}(r)\,h(w)$ in $[0,R)$, $w(0)=\delta$, and blows up at $r=R$. By the strong maximum principle applied to u+vwu+v-w, if $u(0)+v(0)>w(0)$, then $u+v>w$ near $r=0$. However, since $w$ blows up at $R$ while $u+v$ is $C^{2}$ on $[0,R)$, the comparison principle forces $u+v\leq w$ in $[0,R)$ a contradiction with $u(0)+v(0)>w(0)$.
Therefore, $\mathcal{G}_{\delta}$ cannot be unbounded; it is bounded. ∎
Part II: Closedness of $\mathcal{G}_{\delta}$
Proof.
Step 1: Sequential closure argument. Let $\{(\alpha_{m},\beta_{m})\}\subset\mathcal{G}_{\delta}$ be a sequence converging to $(\alpha,\beta)\in\overline{\mathcal{G}_{\delta}}$. For each $m$, Theorem 2.1 guarantees the existence of a positive entire solution $(u_{m},v_{m})$ of (3.1)–(3.2) with
@@
Step 2: Uniform bounds and equicontinuity. From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\alpha_{m}+\beta_{m}$, we deduce that the family $\{(u_{m},v_{m})\}$ is uniformly bounded on compact intervals and equicontinuous.
Step 3: Compactness and passage to the limit. By the Arzelá –Ascoli theorem, there exists a subsequence converging uniformly on compact sets to a limit $(u,v)$. Passing to the limit in (3.1)–(3.2) (justified by the dominated convergence theorem, since the integrands are continuous and uniformly bounded on compacts) shows that $(u,v)$ is a positive entire solution with
@@
Since $\min\{\alpha,\beta\}\geq\delta$ by construction, we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$.
Conclusion: $\mathcal{G}_{\delta}$ is both bounded and closed. ∎
@@
For clarity, we structure the proof of Theorem 2.3 into seven distinct and logically self-contained steps.
Standing assumptions and notation.
@@
We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
@@
We introduce the Keller–Osserman transform
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g(t,t)},\quad s>0.$
By construction:
  • $H$ is strictly decreasing, with
    $H^{\prime}(s)=-\frac{1}{f(s,s)+g(s,s)}<0;$
  • $H^{\prime\prime}(s)\geq 0$ since $s\mapsto f(s,s)+g(s,s)$ is nondecreasing by (H2).
Step 1: Perturbation of initial data and blow-up for approximants
Proof.
From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
Let $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. For each $k\geq 1$, consider the perturbed initial data
@@
If this pair belonged to $\mathcal{G}$, then by downward closure and the definition of $\mathcal{G}_{\delta}$, $(\alpha,\beta)$ would lie in the interior of $\mathcal{G}_{\delta}$, contradicting $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. Hence
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right)\notin\mathcal{G}.$
@@
Step 2: Reduction to a scalar inequality
Proof.
Let
@@
where $(u,v)$ is the limit solution corresponding to $(\alpha,\beta)$. Since $u_{k}\leq W_{k}$ and $v_{k}\leq W_{k}$, and $f,g$ are nondecreasing, we have
@@
$\leq(p(r)+q(r))\,\big(f(W_{k},W_{k})+g(W_{k},W_{k})\big),$ (3.5)
by (H2). ∎
Step 3: Applying the Keller–Osserman transform
Proof.
Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
@@
Since $H^{\prime}(s)=-1/(f(s,s)+g(s,s))$, inequality (3.5) yields
@@
Step 4: First radial integration
Proof.
In radial coordinates, the above inequality becomes
@@
Step 5: Second radial integration and lower bound
Proof.
Integrating from $r$ to $R_{k}$ gives
@@
Letting $r\rightarrow R_{k}^{-}$, we have
$W_{k}(R_{k})\rightarrow\infty\text{ and hence }H(W_{k}(R_{k}))\rightarrow 0.$
Thus
@@
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2). ∎
Step 6: Divergence of at least one component.
Proof.
From the previous steps (in particular, Step 1 of the proof of Theorem 2.2 and (3.4)), we already know that
@@
Suppose, for contradiction, that both $u$ and $v$ are bounded on $0,\infty)$. That is, there exist constants $M_{1},M_{2}>0$ such that
$u(r)\leq M_{1},\quad v(r)\leq M_{2}\quad\forall r\geq 0.$
Then
@@
which contradicts $W(r)\rightarrow+\infty$. Therefore, at least one of the two components must diverge:
@@
This next figure shows
$u\left(r\right)\text{ and }v\left(r\right)\text{ vs. }r,$
with the “both bounded” scenario impossible, and the Keller–Osserman case forcing both curves upward.
[…]
Figure 2: $W(r)\rightarrow+\infty$ forces at least one component to diverge; under the Keller–Osserman condition both diverge. Solid lines: without KO; dashed lines: with KO.
Both components diverge under the supplementary Keller–Osserman condition. Assume now that the additional Keller–Osserman condition at infinity (2.3) holds. We claim that in this case $\lim_{r\rightarrow\infty}u(r)\;=\;\lim_{r\rightarrow\infty}v(r)\;=\;+\infty$. Indeed, if $u$ were bounded, then $v=W-u\rightarrow+\infty$. Thus it suffices to prove the implications
@@
Proof of Claim 1. Assume by contradiction that $u$ is bounded: there exists $U>0$ such that
$0<u(r)\leq U\quad\forall r\geq 0.$
Since $v(r)\rightarrow+\infty$ while $u$ is bounded, there exists $R_{0}>0$ such that
$u(r)\leq U<v(r),\forall r\geq R_{0}.$
By the monotonicity of $f$ in each variable (assumption (H2)), we have
$f\big(u(r),v(r)\big)>f\big(u(r),u(r)\big)=:\ell_{f}\big(u(r)\big)\quad\forall r\geq R_{0},$
where $\ell_{f}(s):=f(s,s)$ is continuous, nondecreasing, satisfies $\ell_{f}(0)=0$, and, by hypothesis, the Keller–Osserman condition
@@
$\Delta u(r)>p(r)\,\ell_{f}\big(u(r)\big).$ (3.6)
Barrier construction. Fix $R_{0}>0$ as above and consider the radial Cauchy problem
$y^{\prime\prime}(r)+\frac{n-1}{r}y^{\prime}(r)$ $=p(r)\,\ell_{f}\big(y(r)\big),\quad r\in[R_{0},R_{a}),$ (3.7)